Provide an appropriate response 8) An Apple Pie Company knows that the number of pies sold each day varies from day to day. The owner believes that on 50% of the days she sells 100 pies. On another 25% of the days she sells 150 pies, and she sells 200 pies on the remaining 25% of the days. To make sure she has enough product, the owner bakes 200 pies each day at a cost of $2 each. Assume any pies that go unsold are thrown out at the end of the day. If she sells the pies for $4 each, find the probability distribution for her daily profit. A) B) Profit P(profit) $200 0.5 $300 0.25 $400 0.25 Profit P(profit) $200 05 $400 0.25 $600 0.25 C) D) Prof P(profit) $0 0.5 $200 0.25 $400 0.25 Profit P(profit) 8400 05 $600 0.25 $800 0.25
Answer #1

Answer

The Correct answer is option (A) i.e

Profit(X) P(Profit)
$100×2 =$200 0.5
$150×2=$300 0.25
$200×2=$400 0.25

As because,

Profit =$4-$2

=$2

Note:-

If you like it then please don't forget to like it.

Thank you

Have a nice day!